please help me solve this word question. The sum of the page numbers on facing pages of a book is 157; what are the page numbers?

Thanks!

Answer by jim_thompson5910(35256)

x+(x+1) = 157

2x + 1 = 157

2x = 157-1

2x = 156

x = 156/2

x = 78

x+1 = 79

So the two page numbers are 78 and 79.